### v1

# On Pretraining Data Diversity for Self-Supervised Learning

Abstract
We explore the impact of training with more diverse datasets, characterized by the number of unique samples, on the performance of self-supervised learning (SSL) under a fixed computational budget. Our findings consistently demonstrate that increasing pretraining data diversity enhances SSL performance, albeit only when the distribution distance to the downstream data is minimal. Notably, even with an exceptionally large pretraining data diversity achieved through methods like web crawling or diffusion-generated data, among other ways, the distribution shift remains a challenge. Our experiments are comprehensive with seven SSL methods using large-scale datasets such as ImageNet and YFCC100M amounting to over 200 GPU days. Code and trained models will be available at https://github.com/hammoudhasan/DiversitySSL.
Keywords:
self-supervised learning distribution shift data diversity
Figure 1: Impact of Diversity on Pretraining: Self-supervised learning (SSL) can be used to pretrain vision models on smaller datasets closely aligned to downstream task data, e.g., pets classification, hence with a small distribution shift (top, wild animals pretraining). Conversely, we could pretrain on an extensively varied dataset, with wide distribution differences (outdoor scenes, bottom). We demistify the role of pretraining diversity in SSL under a fixed computational budget, and highlight its effects in relationship to the distribution shift.
1 Introduction
Self-supervised learning (SSL) has recently emerged as a new paradigm to pretrain large vision models at scale [25, 46, 26]. Leveraging the ability to learn from unlabelled data, pretraining on millions—or even billions [52, 26]—of images turned from an unachievable goal to a common practice. This exposure to extremely diverse datasets, i.e., composed of a remarkable number of unique samples, granted impressive performance and unprecedented generalization capabilities to a growing number of vision models [46]. Large-scale datasets, in conjunction with substantial computational resources, have been the key driving forces for the success of SSL-based approaches. For instance, SEER [25], pretrained for approximately 11 GPU years on a billion images, exemplifies the massive computation and data resources employed for these models. It has become the implicit norm that increasing computation and data is beneficial, without any detailed analysis of how they separately impact SSL effectiveness. In particular, it is not clear to what extent large datasets are responsible for the impressive generalization capabilities of SSL models. Indeed, consider the example in Figure 1. Assuming a fixed monetary budget, in the form of computational expenses, for pretraining a vision model under the SSL paradigm, does iterating over a large set of images work best given a downstream task, or it is better to train repeatedly on a smaller set of samples visually close to the ones of the downstream? In this context, we face the ever-lasting problem of distribution shift on vision models [58]. Without a proper understanding of the impact of data, there could be a wrong allocation of efforts in increasing data or computation, leading to inefficiencies in the deployment process of SSL models.
In this paper, we study the role of the diversity of the SSL pretraining data, under a fixed computational budget, when the pretraining data matches the distribution of the downstream data, as well as when they differ. Our experiments span various SSL methods [10, 14, 4, 9, 28, 65, 8] and multiple datasets [38, 39, 56, 18, 54, 37, 43, 47, 69, 5], and are tested under several computational budgets amounting to a total of 200 GPU days. We summarize our contributions below:
  1. 1.
    We show that SSL pretraining strategies are currently data-inefficient in compensating for distribution shifts. Under normalized computational costs, we verify that pretraining on large datasets with high diversity cannot outperform the pretraining on datasets with limited diversity, but with a lower distribution shift with respect to the downstream tasks.
  2. 2.
    We conclude that there is a wide margin for improvement in the performance of current SSL methods on out-of-distribution classes, and we propose insights for a fair evaluation. This shows the need for future research on new SSL techniques leveraging data diversity better, to improve generalization capabilities beyond the training data distribution.
  3. 3.
    We propose a novel strategy for a computationally-normalized evaluation of SSL and a carefully designed set of experiments focusing on pretraining data diversity, enabling us to draw our novel conclusions.
Ultimately, our work provides a comprehensive analysis, from experiments worth 200 GPU days, on the interaction between data diversity and computational resources, and their impact on the performance of SSL models, with an aim to improve pretraining practices. Now, we will analyze the relevant literature.
2 Related Work
Self-Supervised Learning   Early work in self-supervised learning used simple pretext tasks, such as relative patch prediction [19, 20], image colorization [67], image rotation prediction [24], or solving jigsaw puzzles [44] to train feature extractors in absence of annotations. Representations learned with those methods are limitedly effective, hence more recent literature has moved towards more sophisticated approaches, such as using image augmentations to generate correlated views of a training sample, and learning to extract augmentation-invariant representations for these correlated pairs. Among these multi-view methods, many exploit contrastive losses [8, 10, 45, 35, 11, 12, 31, 41, 1, 34, 42], enforcing similarity between views of the same image (positives) and dissimilarity from other images (negatives). Due to the need of many negative samples, contrastive methods often require large batch sizes to work effectively [10, 14]. Cluster-based methods such as SwAV [8], DINO [9] and DeepCluster v2 [6] learn generalizable representations by grouping samples into cluster prototypes. Others exploit prediction of features with siamese networks [13], learn features in a teacher-student fashion [28], or use redundancy reduction techniques [65, 4]. Finally, masked image modeling [3, 30, 63] emerged as a scalable alternative for Vision Transformers [21], learning representations by predicting masked image patches as SSL task.
Pretraining at Scale   SSL is most effective when pretraining is conducted at scale, benefiting from large datasets and great computational resources. Initial attempts at large-scale pretraining were made via combining contrastive learning and clustering [55, 7]. The SEER model [25] was trained on 1 billion internet images using SwAV [8] applied on a RegNet [48] backbone. The importance of model scaling to leverage large-scale datasets is also shown in  [64, 27, 26], as well as the need for increasing training duration [64]. Additional strategies are necessary to achieve best performance at scale [66, 17, 23]. All considered works focus on reaching the best representations, without many considerations about training costs, thus encouraging unfair comparisons about the efficacy of data. A preliminary work [15] found for SimCLR [10] that increasing data from 500 thousand to 1 million images leads to a modest boost in downstream performance, but without limiting the training budget. While large uncurated datasets have been further explored [55], their efficacy in relationship with the distribution shift under normalized computation is still limitedly investigated.
Distribution Shift in SSL   Some studies have investigated how the pretraining data domain affects downstream performance on other domains. For various pretraining datasets, preliminary works [36, 15, 27, 40] observed that the best performing representations were learned on datasets that were similar to the downstream test task. Additionally, combining datasets before pretraining or combining self-supervised features learned from various datasets did not lead to significant improvements in [15]. In  [68, 51], they showcase higher generalization capabilities of SSL models compared to their supervised counterparts, for several downstream tasks in the presence of a distribution shift. In [59], they pretrained on several datasets, and observed different generalization depending on the object-centric or scene-centric appearance of the pretraining dataset. Furthermore, initial considerations on the impact of external data on downstream tasks under distribution shift have been proposed in [60]. Some compensate the distribution shift with the scale of training datasets [29]. Although partial information can be inferred by these works, there is still a lack of a fair, computation-normalized evaluation that allows to study the effects of the distribution shift in a controlled environment.
3 Preliminaries
Pretraining   We first outline the general pretraining procedure common to state-of-the-art self-supervised learning methods. Specific differences among these methods are detailed in the supplementary material. The overall pretraining pipeline, common across many SSL approaches [10, 28, 65, 9, 8, 4], goes as follows: (1) Data Sampling: from a large-scale, unlabeled upstream pretraining dataset, denoted as $\mathbb{D}_{\text{SSL}}$, an image $\mathbf{x}$ is randomly sampled; (2) View Generation: two correlated views, $\tilde{\mathbf{x}}_{A}$ and $\tilde{\mathbf{x}}_{B}$, are generated from $\mathbf{x}$ using two image augmentation functions sampled at random. These random augmentations include random resized cropping, horizontal flipping, blurring, and color adjustments [2], among others; (3) Feature Extraction and Projection: the correlated views undergo feature extraction through a network, $f_{\theta_{f}}$, parameterized by $\theta_{f}$, such as a ResNet [32] or a ViT [21], leading to representations $\mathbf{h}_{A}=f_{\theta_{f}}(\tilde{\mathbf{x}}_{A})$ and $\mathbf{h}_{B}=f_{\theta_{f}}(\tilde{\mathbf{x}}_{B})$. A linear projection head, $g_{\theta_{g}}$, parameterized by $\theta_{g}$, then maps these representations to a latent space, resulting in $\mathbf{z}_{A}=g_{\theta_{g}}(\mathbf{h}_{A})$ and $\mathbf{z}_{B}=g_{\theta_{g}}(\mathbf{h}_{B})$; (4) Joint Optimization: The feature extractor $f_{\theta_{f}}$ and the projection head $g_{\theta_{g}}$ are jointly optimized according to the following objective:
$\theta_{f}^{*},\theta_{g}^{*}=\underset{\theta_{f},\theta_{g}}{\arg\min}~{}\mathbb{E}_{\mathbf{x}\sim\mathbb{D}_{\text{SSL}}}~{}\mathcal{L}_{\text{SSL}}(\mathbf{z}_{A},\mathbf{z}_{B}),$ (1)
where $\mathcal{L}_{\text{SSL}}$ is a loss function specific to the chosen SSL pretraining method.
After pretraining, the feature extractor $f_{\theta_{f}^{*}}$ can be deployed for various downstream tasks such as image classification, object detection, or segmentation. This is typically achieved by training a task-specific head. Alternatively, the feature extractor $f_{\theta_{f}^{*}}$ can either be fine-tuned or used together with a $k$-nearest neighbors (kNN) classifier.
Linear Probing   There are several ways to evaluate the performance of a self-supervised learning method such as linear probing [28, 14, 10], kNN [62, 71, 9, 8], and few-shot evaluation [25, 22]. Consistent with the general protocol established in the literature [10, 31, 67, 15], we use linear probing to measure the quality of the features extracted for classification tasks. The procedure is as follows: (1) a labeled downstream dataset, $\mathbb{D}_{\text{task}}$, consisting of image-class pairs $(\mathbf{x},\mathbf{y})\sim\mathbb{D}_{\text{task}}$, is selected for evaluation. (2) For each image $\mathbf{x}$, its representation is extracted using the pretrained feature extractor $f_{\theta_{f}^{*}}$, after which the linear classification head $t_{\theta_{t}}$, parameterized by $\theta_{t}$, is then applied to obtain $t_{\theta_{t}}(f_{\theta_{f}^{*}}(\mathbf{x}))$. (3) The linear head $t_{\theta_{t}}$ is optimized as follows:
$\theta^{*}_{t}={\underset{\theta_{t}}{\arg\min}~{}\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim\mathbb{D}_{\text{task}}}\left[\mathcal{L}_{\text{task}}(t_{\theta_{t}}(f_{\theta_{f}^{*}}(\mathbf{x})),\mathbf{y})\right]},$ (2)
Note that only the parameters of the linear head $\theta_{t}$ are optimized, while the parameters $\theta^{*}_{f}$ of the feature extractor are kept frozen. The quality of features extracted by $f_{\theta_{f}^{*}}$ is directly inferred from the classification accuracy achieved on the test set of $\mathbb{D}_{\text{task}}$, which serves as the primary indicator of the quality of the extracted features.
4 Normalized Evaluation
We stress how for a correct understanding of the impact of data diversity we need to analyze its effects isolating them from the impact of increased computation. To enable this, we introduce (1) a computational budget used to normalize computation across experiments, and (2) a quantification of the data diversity seen by models during pretraining.
Computational Budget   Current progress in SSL pretraining simultaneously scales computational budget and dataset size to achieve the best performance [25]. This makes it difficult to assess the reasons behind the success of many SSL algorithms. Do these methods benefit mainly from the large amounts of computation, i.e., running SSL pretraining for large numbers of epochs, or do they benefit from data diversity in larger datasets containing a vast variety of visual features? To perform a disentangled evaluation of these two factors, we first introduce $\mathcal{C}$ as a measure of computational budget, which quantifies the total number of images an SSL method is allowed to process during pretraining. This is calculated as $\mathcal{C}=N\cdot\mathcal{E}$, where $N$ is the number of unique samples in the pretraining dataset $\mathbb{D}_{\text{SSL}}$, hence the data diversity of the dataset, and $\mathcal{E}$ is the number of pretraining epochs. Constraining multiple models pretrained with SSL to a fixed computational budget $\mathcal{C}$ allows for meaningful comparison, as it is guaranteed that all SSL methods will have processed the same number of pretraining images.
Table 1: Impact of Data Diversity on CIFAR-100 and Tiny ImageNet SSL Pretraining Performance: We study the effects of diversity on CIFAR-100 1(a) and Tiny ImageNet 1(b) across seven different methods and three data diversity settings for a ResNet-18 pretraining, where for all, $\mathbb{D}_{\text{task}}=\mathbb{D}_{\text{SSL}}$. This comparison includes analyzing classification accuracy through linear probing on the training set and evaluation on the test set of the respective datasets. Although performance fluctuates among different methods, a consistent trend is observed: higher data diversity typically leads to the generation of higher quality representations.
$N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
(×$(\times 10^{3})$ (×$(\times 10^{-3})$ SimCLR B.T. BYOL SwAV VICReg MoCoV3 DINO
5 0.1
37.95
44.30
41.14
14.06
44.06
12.51
15.74
25 0.5
50.13
54.29
51.01
45.44
50.37
51.58
35.89
50 1.0 58.59 58.42 58.28 56.37 55.83 55.61 40.93
$\mathcal{C}=50\times 10^{6}$
(a) CIFAR-100
$N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
(×$(\times 10^{3})$ (×$(\times 10^{-3})$ SimCLR B.T. BYOL SwAV VICReg MoCoV3 DINO
10 0.2
36.91
40.98
35.62
34.23
37.58
36.56
34.54
50 1.0
48.77
52.01
48.05
43.63
48.75
46.45
44.39
100 2.0 49.83 55.60 50.29 47.48 54.10 48.58 47.32
$\mathcal{C}=50\times 10^{6}$
(b) Tiny ImageNet
Quantifying Pretraining Diversity   Even under normalized computation $\mathcal{C}$, various SSL approaches may be exposed to different data diversity as they are trained with different datasets. For instance, a model trained for $\mathcal{E}=1000$ epochs on a dataset of size $N=1000$ will see less diversity than a model pretrained for $\mathcal{E}=1$ epoch on a dataset of size $N=10^{6}$, despite that they are both pretrained under normalized computation, processing the same number of images. Hence, to capture this notion of exposure to diversity while pretraining under normalized computation $\mathcal{C}$, we define a pretraining diversity $\mathcal{D}$, which captures the number of unique samples encountered during training given a fixed $\mathcal{C}$ as $\mathcal{D}=\nicefrac{{N}}{{\mathcal{C}}}=\nicefrac{{1}}{{\mathcal{E}}}$. A model pretrained with larger $\mathcal{D}$ indicates that the model is presented with a larger number of unique images during training with fixed $\mathcal{C}$, and hence is exposed to more pretraining data diversity. In the next section, we explore the effects of variations in $\mathcal{D}$ on SSL performance under a distribution shift, while keeping $\mathcal{C}$ constant.
5 Fixed Budget SSL: In & Out-of-Distribution
Training Configuration   We evaluate seven SSL methods: SimCLR [10], MoCoV3 [14], VICReg [4], DINO [9], BYOL [28], Barlow Twins [65], and SwAV [8]. We use different datasets for both pretraining $\mathbb{D}_{\text{SSL}}$ and linear probing $\mathbb{D}_{\text{task}}$ for different sections. We use solo-learn [16] as a codebase. For each method, we use the default parameters provided when available, otherwise, we conduct a hyperparameter search to ensure proper convergence.
5.1 Performance on the Same Distribution
We aim to answer a first question: does collecting more samples from the same distribution help SSL pretraining with a fixed budget? We conduct experiments to capture how the pretraining diversity $\mathcal{D}$ influences SSL pretraining within a normalized $\mathcal{C}$, focusing on the simplest setting where the upstream and downstream datasets belong to the same distribution such that $\mathbb{D}_{\text{SSL}}=\mathbb{D}_{\text{task}}$. This serves as a fundamental ground for our further experiments.
Setup   For pretraining, we use CIFAR-100 [38] and Tiny ImageNet [39] as $\mathbb{D}_{\text{SSL}}$, which contain $50\times 10^{3}$ and $100\times 10^{3}$ images, respectively. We set $\mathcal{C}=50\times 10^{6}$, chosen such that it allows all methods to converge during pretraining. In Section 6, we study the effect of varying the budget $\mathcal{C}$. We pretrain on subsets of $\mathbb{D}_{\text{SSL}}$ with different sizes $N$ (10%, 50%, and 100% of $\mathbb{D}_{\text{SSL}}$), enabling us to observe the effects of pretraining diversity $\mathcal{D}$ on training outcomes where $\mathcal{E}$ is adjusted accordingly to match the budget $\mathcal{C}$. For example, using 100% of CIFAR-100 involves 1000 epochs of training, while 10% and 50% of the dataset lead to 10000 and 2000 epochs, respectively. These subsets of $\mathbb{D}_{\text{SSL}}$ are created by shuffling the dataset and then selecting the first $10\%,50\%$, or $100\%$ split. All models in this section use a ResNet-18 [32] backbone pretrained from scratch. For evaluation, we use in-distribution linear probing, i.e., $\mathbb{D}_{\text{task}}=\mathbb{D}_{\text{SSL}}$, where performance is measured by classification accuracy on the test set.
Results   The results of linear probing are presented in Tables 1(a) and 1(b) for CIFAR-100 and Tiny ImageNet, respectively. While different SSL methods show varying levels of performance, we observe that, for all methods, an increase in the diversity $\mathcal{D}$ consistently leads to higher linear probing accuracy, suggesting that including more in-distribution samples in $\mathbb{D}_{\text{SSL}}$ helps under a normalized $\mathcal{C}$. For example, SimCLR achieves an accuracy of 37.95 when 10% of CIFAR-100 is provided, whereas this performance improve by around 12% when only 50% of the dataset is provided and by another 8% after pretraining on 100% of unique samples. This suggests that SSL methods substantially benefit from having a more diverse pretraining set in computationally budgeted in-distribution evaluation, a fundamental verification that allows us to proceed in our analysis.
Insight (1)
When the distributions of the upstream and downstream tasks are the same, i.e., $\mathbb{D}_{\text{SSL}}$ = $\mathbb{D}_{\text{task}}$, in a normalized computation setup, increasing pretraining diversity $\mathcal{D}$ proves to be an effective strategy for enhancing SSL pretraining performance.
5.2 Increasing Data Diversity
As observed in Section 5.1, if $\mathbb{D}_{\text{SSL}}=\mathbb{D}_{\text{task}}$ having a more diverse pretraining set benefits SSL methods, under a normalized computation assumption. However, to increase diversity, sampling from the same distribution $\mathbb{D}_{\text{task}}$ to extend $\mathbb{D}_{\text{SSL}}$ is not always attainable. In fact, pretraining data is often sampled from a distribution different than $\mathbb{D}_{\text{task}}$. The added samples will then introduce a distribution shift between $\mathbb{D}_{\text{SSL}}$ and $\mathbb{D}_{\text{task}}$.
[…]
Figure 2: Data Collection Strategies: We analyze strategies for collecting additional data ($\mathbb{A}$), i.e., collecting more source data, crawling the web or using synthetic images. Using a class prior (top row) simulates In-distribution trainings. We also collect images without class prior (bottom row) to analize the interactions between diversity and Out-of-distribution classes.
Setup   We use ImageNet-100 [56] as $\mathbb{D}_{\text{task}}$, and we construct multiple $\mathbb{D}_{\text{SSL}}$ to evaluate the effects of different data collection strategies and the distribution shift. We first introduce a set $\mathbb{B}$ composed of $65\times 10^{3}$ images from ImageNet-100 (50% of the dataset), which we use as a baseline for $\mathbb{D}_{\text{SSL}}$ with minimum diversity. We denote the 100 classes in ImageNet-100 as $\mathcal{T}_{\text{100}}$, and sample $\mathbb{B}$ uniformly including images from all classes. Next, we compare with pretraining on more diverse datasets as $\mathbb{D}_{\text{SSL}}$, imposing $\mathbb{D}_{\text{SSL}}=\mathbb{B}\cup\mathbb{A}$ where $\mathbb{A}$ includes $65\times 10^{3}$ images sampled with one of three strategies. To highlight the effects of the distribution shift, we include in $\mathbb{A}$ either images from In-distribution classes, i.e., selecting images from classes overlapping with $\mathcal{T}_{\text{100}}$, or images from Out-of-distribution classes, which do not overlap with $\mathcal{T}_{\text{100}}$. This is to study the effects of the distribution shift, since we do not assume access to downstream classes in real scenarios. For the Out-of-distribution samples, we define $\mathbb{A}$ as (1) random images sampled from the full ImageNet [18] dataset; (2) images crawled from Flickr, Bing, and DuckDuckGo; or (3) images generated with Stable Diffusion V2.1 [50]. We respectively call these sets $\mathbb{A}_{\text{Source}}^{\text{{{Out}}}}$, $\mathbb{A}_{\text{Web}}^{\text{{{Out}}}}$, and $\mathbb{A}_{\text{Synthetic}}^{\text{{{Out}}}}$. We also define their In-distribution counterparts as $\mathbb{A}_{\text{Source}}^{\text{{{In}}}}$, $\mathbb{A}_{\text{Web}}^{\text{{{In}}}}$, and $\mathbb{A}_{\text{Synthetic}}^{\text{{{In}}}}$, respectively. Note how $\mathbb{B}\cup{\color[rgb]{0.87890625,0.4375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.87890625,0.4375,0}\pgfsys@color@rgb@stroke{0.87890625}{0.4375}{0}\pgfsys@color@rgb@fill{0.87890625}{0.4375}{0}\mathbb{A}_{\text{Source}}^{\text{{{In}}}}}$ is the full ImageNet-100, coherently with Section 5.1. Figure 2 shows each collection strategy, for which we provide implementation details in supplementary material. Although many factors (such as the appearance of images) impact the distribution shift, using a class prior imposes that any strategy using it would still result in closer distribution with respect to the same strategy without class priors [15]. We pretrain a ResNet-18 from scratch, with the same settings of Section 5.1, and $\mathcal{C}=50\times 10^{6}$. Note that this results in $\mathcal{D}=1.25\times 10^{-3}$ for pretraining on $\mathbb{B}$, and $\mathcal{D}=2.5\times 10^{-3}$ for pretraining on $\mathbb{B}\cup\mathbb{A}$, introducing a $\mathcal{D}$ difference of a factor of 2.
Figure 3: Effect of Various Data Sources on SSL Pretraining: We use a baseline set $\mathbb{B}$ (black dashed line), comprising $65\times 10^{3}$ images from ImageNet-100, for pretraining a ResNet-18 with $\mathcal{C}=50\times 10^{6}$. Augmenting $\mathbb{B}$ with In-distribution images enhances performance (above black line), while Out-of-distribution augmentations reduce it (below black line).
Results   We report the linear probing accuracies for pretraining on each $\mathbb{B}\cup\mathbb{A}$ as bars in Figure 3, showing the $\mathcal{C}$-normalized training on $\mathbb{B}$ as a dashed line. Surprisingly, without class priors (Out), including $\mathbb{A}_{\text{Source}}^{\text{{{Out}}}}$, $\mathbb{A}_{\text{Web}}^{\text{{{Out}}}}$, and $\mathbb{A}_{\text{Synthetic}}^{\text{{{Out}}}}$ underperforms compared to pretraining on $\mathbb{B}$ only. For instance, for SimCLR and while $\mathbb{B}$ scores 72.3% accuracy, increasing diversity reduces the accuracy by 1% in all cases. This might appear to conflict with our findings in Section 5.1, however, the inclusion of Out samples leads to $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$, since we sample only classes not included in $\mathbb{D}_{\text{task}}$. We infer that, with normalized $\mathcal{C}$, increasing $\mathcal{D}$ without distribution priors may negatively impact performance. Conversely, when class priors are available (In), increasing pretraining diversity $\mathcal{D}$ improves performance compared to $\mathbb{B}$ pretraining. For instance, pretraining on $\mathbb{A}_{\text{Web}}^{\text{{{In}}}}$ performs comparably to augmenting with additional ImageNet samples ($\mathbb{A}_{\text{Source}}^{\text{{{In}}}}$), as in the case of SimCLR where the inclusion of $\mathbb{A}_{\text{Web}}^{\text{{{In}}}}$ scores only 0.2% lower than $\mathbb{A}_{\text{Source}}^{\text{{{In}}}}$. Including $\mathbb{A}_{\text{Synthetic}}^{\text{{{In}}}}$ data also helps, Synthetic $\mathbb{A}_{\text{Synthetic}}^{\text{{{In}}}}$ data helps, although more limitedly due to the visual discrepancies between generated and real images. Ultimately, the effectiveness of diversity is linked to the distribution shift. These findings highlight the impact of the distribution shift on computationally-normalized SSL pretraining and help define evaluation practices for this task (see Section 7).
Insight (2)
When the distributions of the upstream and downstream tasks differ, $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$, and in a normalized computation setup, increasing pretraining diversity $\mathcal{D}$ may harm SSL pretraining performance, reflecting the adverse effects of distribution shift.
Table 2: Performance of ImageNet and YFCC100M SSL Pretraining on Various Downstream Tasks: We train ResNet-50 and ViT-B/16 under normalized computation ($\mathcal{C}=98\times 10^{6}$) using SimCLR (left) and MoCoV3 (right) on ImageNet and YFCC100M (YFCC) with multiple $\mathcal{D}$. Despite the significantly larger $\mathcal{D}$ when trained on YFCC100M, these models cannot offset the effects of the distribution shift and are outperformed by models pretrained on ImageNet in the majority of downstreams.
[…]
SimCLR
$\mathbb{D}_{\text{SSL}}$ $N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
(×$(\times 10^{6})$ (×$(\times 10^{-3})$ ImageNet Cars Flow. Pets Places Food
ResNet-50 ImageNet 0.128 1.31 56.9 43.0 82.3 73.9 45.4 59.9
ImageNet 0.256 2.61 61.1 45.5 84.0 76.0 47.0 64.4
ImageNet 0.640 6.54 63.7 46.8 84.2 78.8 48.3 67.2
ImageNet 1.281 13.0 64.5 46.4 84.6 79.5 48.8 68.0
[…]
ViT-B/16 ImageNet 0.128 1.31 54.2 37.2 81.8 70.3 44.6 64.3
ImageNet 0.256 2.61 61.3 39.5 83.8 77.7 47.2 69.4
ImageNet 0.640 6.54 65.5 39.9 84.6 80.4 49.1 73.1
ImageNet 1.281 13.0 66.7 39.5 83.7 81.7 50.0 73.6
[…]
$\mathcal{C}=98\times 10^{6}$
MoCoV3
$\mathbb{D}_{\text{SSL}}$ $N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
(×$(\times 10^{6})$ (×$(\times 10^{-3})$ ImageNet Cars Flow. Pets Places Food
ResNet-50 ImageNet 0.128 1.31 58.1 40.6 81.8 76.04 45.1 63.9
ImageNet 0.256 2.61 62.9 45.3 85.0 79.8 47.6 68.7
ImageNet 0.640 6.54 65.4 48.4 86.1 81.9 49.2 71.0
ImageNet 1.281 13.0 65.9 48.8 86.6 82.6 49.5 71.9
[…]
ViT-B/16 ImageNet 0.128 1.31 57.9 33.4 82.8 78.0 46.7 67.8
ImageNet 0.256 2.61 63.7 35.0 85.3 82.9 48.4 71.0
ImageNet 0.640 6.54 67.2 39.5 85.0 85.8 49.7 72.8
ImageNet 1.281 13.0 68.8 41.9 86.5 86.5 50.3 73.8
[…]
$\mathcal{C}=98\times 10^{6}$
Model Saturation   In Section 5.3, we evaluated the effects of extreme differences in $\mathcal{D}$. We aim to understand the trend in pretraining on YFCC100M, and if adding more samples could compensate the distribution shift with ImageNet. Hence, we extend the YFCC100M experiments from Section 5.3, examining various subsets—specifically 0.1%, 1%, 10%, and 100% of the dataset. Again, we normalize the computational budget to $\mathcal{C}=98\times 10^{6}$, equivalent to one epoch on YFCC100M. The results of linear probing models pretrained using SimCLR and MoCoV3 with ResNet-50 and ViT-B/16 on ImageNet are shown in Figure 5.2. Interestingly, we obtain a performance plateau. This observation points to a saturation point in performance, showing that simply increasing the dataset further would not bridge the gap between the $\mathbb{D}_{\text{SSL}}$ and $\mathbb{D}_{\text{task}}$. Hence, arbitrarily increasing pretraining diversity $\mathcal{D}$ is not sufficient to bridge the distribution shift problem. Here $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$, hence still aligning with our findings in Section  5.1.
Label Quantity   A question that arises in our setting is whether having a higher pretraining diversity leads to requiring less labeled samples during linear probing. In this section, we focus on understanding the impact of increased $\mathcal{D}$ on the number of labeled samples required for effective linear probing in downstream tasks. We use the same trained models from Section 5.3, i.e., SimCLR and MoCoV3, with ResNet-50 and ViT-B/16 architectures. In this setup, we set $\mathbb{D}_{\text{task}}=\text{ImageNet}$, with two upstream dataset scenarios: ImageNet (where $\mathbb{D}_{\text{SSL}}=\mathbb{D}_{\text{task}}$) and YFCC100M (where $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$). Our experiments are summarized in Table 5. We note that with ViT-B/16, if ImageNet is used for pretraining, linear probing with just 5% labeled samples can surpass the performance achieved using 100% labeled data when YFCC100M serves as $\mathbb{D}_{\text{SSL}}$. This also implies that when $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$, a higher quantity of labeled data is necessary to attain competitive performance, and increasing $\mathcal{D}$ does not bring considerable benefits on label quantity under distribution shift. This implies that our findings in Section 5.3 hold regardless of the linear probing labeled set. We note that in scenarios where $\mathbb{D}_{\text{SSL}}=\mathbb{D}_{\text{task}}$, using only 50% of the labeled data can achieve similar performance as utilizing the full 100% of labeled samples, implying that increasing $\mathcal{D}$ leads to reduced label requirement efforts for downstream tasks.
Table 5: Evaluating Network Accuracy With Varied Label Quantity: We evaluate the accuracy of networks trained on ImageNet and YFCC with different labeling percentages of $\mathbb{D}_{\text{task}}=\text{ImageNet}$. The increased diversity still does not compensate for the distribution shift. However, for in-distribution data, one can get away with fewer labels with more diverse pretraining data.
SimCLR
Network $\mathbb{D}_{\text{SSL}}$ $N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
(×$(\times 10^{6})$ (×$(\times 10^{-3})$ 5% 10% 50% 100%
ResNet-50 ImageNet 1.28 1.31 50.3 54.2 61.7 64.5
YFCC 98.2 1000 39.4 44.3 54.0 57.3
ViT-B/16 ImageNet 1.28 1.31 55.2 59.4 65.6 66.7
YFCC 98.2 1000 40.7 45.7 53.3 54.5
$\mathcal{C}=98\times 10^{6}$
MoCoV3
Network $\mathbb{D}_{\text{SSL}}$ $N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
(×$(\times 10^{6})$ (×$(\times 10^{-3})$ 5% 10% 50% 100%
ResNet-50 ImageNet 1.28 1.31 52.1 56.2 63.8 65.9
YFCC 98.2 1000 42.8 48.2 57.7 60.4
ViT-B/16 ImageNet 1.28 1.31 59.3 63.0 68.3 68.8
YFCC 98.2 1000 43.9 49.0 56.0 57.2
$\mathcal{C}=98\times 10^{6}$
[…]